OpenStudy (anonymous):

Nicole has a clay vase shaped as a triangular prism with a base area of 136.9 square inches. It is one third filled with water. If there are 575 cubic inches of water in the vase, what is the height of the vase, rounded to the nearest tenth? Answer 7.1 inches 8.3 inches 12.6 inches 13.5 inches

OpenStudy (hoblos):

the vase is one third filled with water means that 575 cubic inches is one third of the total volume of the vase so V=1725 cubic inches in general.. V = area of base x height so height = V/(area of base) = 1725/136.9 = 12.6 inches

OpenStudy (anonymous):

thank-you!!!!

OpenStudy (hoblos):

its my pleasure :))

OpenStudy (anonymous):

=)